OpenStudy (anonymous):

Combine the radicals Below -√5-3√5

OpenStudy (callisto):

You may treat √5 as x So, -√5-3√5 = -x -3x = ... (What do you get in terms of x?)

OpenStudy (callisto):

After you find that in terms of x, sub x=√5 back to that term. Then, you can get your answer.